# Jingxuan Li UCLA

### v2

@@ 5 Inducing VLM’s Preferences
Our initial experiment showed that some VLMs have inherent inclinations toward different values. We now explore the dynamic aspects of VLM preferences beyond these static traits. We use Role-Playing Language Agents (RPLA) (Chen et al., 2024b) as a framework to assess VLMs’ ability to adapt dynamically and simulate different personas, making decisions accordingly. Building on research showing that LLMs can emulate personas through RPLA (Serapio-García et al., 2023), we pose two key questions for VLMs: (1) How well can VLMs align their traits to role-play personas using specific prompts? (2) Can strategies enhance accuracy and consistency in role-playing performance?
@@ 5.1 Experiment
We use social media recommendation systems to evaluate whether VLMs can exhibit preferences aligned with the specified embedded persona. These systems rely on viewing duration as a key signal for content recommendation(Appendix  A). We assess the VLMs’ role-playing ability by analyzing how well the recommended content reflects the imposed preferences. For example, adopting a pet owner persona should heighten the model’s emphasis on Benevolence, valuing kindness and care, resulting in longer engagement with pet care videos and increased related recommendations Liu et al. (2023).
In addition, we improve VLM performance on social networks by inducing personas through a questionnaire Abeysinghe and Circi (2024), systematically evaluating traits like emotional engagement, value alignment, curiosity, and preference matching to guide structured optimization.
@@
The VLM engages with video shorts by responding either yes or no. A yes response prompts the VLM to remain on the current video, while a no results in an immediate skip. Alignment is measured as the increase in the frequency of recommended content the VLM decides is interesting over time.
@@
We define $I_{\text{avg}}$, the averaged percentage increase of yes responses, to measure the effectiveness of the strategy. $Y_{l}(i)$ and $Y_{f}(i)$ are the number of yes responses until $i$-th video in the last and first $n=50$ videos, respectively. For each model, we conducted $N=10$ trials, with each trial consisting of $100$ video scrolls in total.
@@
Model-wise, CogVLM and Qwen-VL-Plus viewed all Schwartz values favorably, but CogVLM excelled in role-playing, effectively adopting role-specific preferences, while Qwen-VL-Plus showed only partial adherence. Blip-2 demonstrated no engagement or role-playing ability, lacking any signs of an induced personality. The findings show that even basic prompts can evoke detectable preferences, with some platforms emerging as particularly well-suited for role-playing tasks. Model adaptability in expressing role-related traits varied significantly if the persona was given in a simpler prompt.
Inductive Scoring Questionaire Strategy.
Figure 9: Value Distribution Comparison between VLMs and LLMs. For the same model (e.g., GPT-4o and GPT-4o_text), different input modes (multi-modal vs. text-only) are compared. Experiments demonstrate that the choice of multi-modal input significantly influences some models’ value preferences. While models like GPT-4o show consistency across input modes, others, such as Claude 3.5 Sonnet and Gemini 1.5 Pro, exhibit notable differences in preferences
Building on insights from the simple questioning approach, we developed the Inductive Scoring Questionnaire (ISQ) to enhance VLMs’ performance in social media alignment tasks. ISQ employs a series of prompts inquiring about various aspects of the screenshot. When presented with visual content, VLMs are asked to rate aspects like visual appeal, preference alignment, curiosity, etc.
@@
The ISQ calculates a composite score to assess VLM engagement, with scores above a threshold (e.g. $60$) indicating genuine interest, prompting extended interaction. This layered approach enhances persona analysis and final preference evaluation, improving role-specific performance on social media platforms.
@@
Following the trend in simple strategy, we could see a strong increase for TikTok, particularly with Gemini 1.5 Pro, which demonstrates an average rise of as much as $51.9$. GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet—all displayed notable improvements across platforms, with consistently positive changes in performance. This suggests that these models respond well to the ISQ strategy, allowing them to adopt and express induced personalities with greater depth. Among them, Gemini 1.5 Pro and Claude 3.5 Sonnet particularly benefitted from the ISQ approach, showing remarkable growth in comparison to earlier results in the simple strategy. This demonstrates that the ISQ strategy enhances their ability to engage with role-playing tasks more effectively than the previous simple strategy.
@@ 6 Discussion
LLM vs.VLM. We conducted experiments to examine whether different multi-modal inputs influence model value preference outcomes. As shown in Fig. 9, we compared value preferences derived directly from VLMs using images as input with those generated by feeding the corresponding text-based image descriptions created by the same VLM into their paired LLMs. The results reveal significant differences in value preferences between these two modes. For many value dimensions, VLMs and LLM produced distinct preference distributions, especially in models like Claude 3.5 Sonnet and Gemini 1.5 Pro, where the outputs diverged significantly across input modes. In contrast, GPT-4o displayed greater consistency across modes, suggesting its ability to integrate visual and textual information cohesively. These findings highlight that the choice of input mode—visual or text—can significantly affect model outputs, underscoring the importance of input selection in applications requiring personalized or human-like responses. Detailed evaluation methods and results are provided in Appendix E.
Single Frame Screenshot Representation. To validate single-frame screenshots for video content analysis, we randomly selected 500 images from each of TikTok, Instagram, and YouTube for human evaluation. Annotators were provided with the instructions outlined in Appendix D, along with the images and additional context for their judgments. Each image-video pair was assessed by three annotators, resulting in a total of 4,500 ratings. Annotators reviewed the full video and its corresponding screenshot, rating how accurately the screenshot represented the video’s content. The results indicate that 90.4% of screenshots were deemed representative of the video’s main content, demonstrating the effectiveness of single-frame screenshots across platforms. However, 8.8% of the frames were rated as non-representative, highlighting the challenges posed by videos with complex scenes or rapid transitions.
@@ 7 Conclusion
This study introduced Value-Spectrum, a benchmark for evaluating value preferences in VLMs using a vector database derived from social media platforms. Through systematic evaluation, we observed a shared global inclination among models toward certain mainstream values, such as Hedonism, likely influenced by the nature of their training data. At the same time, significant differences emerged across other value dimensions, highlighting disparities in how VLMs align with diverse human-designed value systems. These findings reveal both commonalities that reflect broader societal trends and divergences that underscore model-specific characteristics, prompting us to explore whether these variations can be systematically adjusted to induce specific personas.
This work provides practical insights into VLMs’ ability to adapt their value preferences dynamically through role-playing, offering a pathway to align machine behaviors with human-designed personas. By connecting role-playing capabilities and alignment strategies, we aim to inspire further research into value-driven AI agent systems and their adaptability in real-world applications.
@@ 8 Limitations
The evaluation utilizes Schwartz value dimensions as the foundation for understanding personality traits and preferences, highlighting opportunities for future research to incorporate broader cultural and personality-based perspectives. Future studies might consider expanding the set of value dimensions or integrating alternative value systems, which could further enrich the understanding of diverse value traits. Additionally, even though our use of single-frame screenshots to represent video content proved effective, human evaluators rated the representativeness highly. This approach simplifies analysis, though it may present challenges for capturing the essence of videos with highly dynamic or complex scenes, offering an area for future refinement.
@@ B.1 Experiment Steps
Open The Designated Social Media Platform The second step involves accessing the designated social media platform. For demonstration, we focus on TikTok and GPT4o.
  • Open TikTok website(fig. 10).
  • Navigate to the ’For You’ page where a variety of content is displayed.
@@
Capture Screenshot Image of Playing Short Video Next, we capture screenshots of the short videos that are playing. This captured screenshot is then input to the VLM. An example of a screenshot is shown in fig. 11 (URL: https://www.tiktok.com/@pugloulou/video/7342967563321822497).
Figure 11: Screenshot of playing short video
@@
Simple Strategy: See details in fig. 13.
ISQ Strategy: See details in fig. 14.
@@
Preference Induction: For recommendation systems to recognize user preferences, staying duration for each video is critical. Capturing multiple frames increases processing time, causing most videos to be viewed in their entirety before scrolling. This diminishes the strategy’s impact and hinders the system’s ability to distinguish preferences between videos.
@@ Appendix D Human Annotators: Single-Frame Analysis
We refine the survey formats provided to annotators through multiple iterations, conducting pilot studies on Amazon Mechanical Turk (MTurk) to continuously adjust the instructions until the quality of answers by the annotators meets the desired standard. The instruction examples referenced by the annotators can be found in Fig. 15.
In Amazon MTurk task description provided to annotators, we clearly stated that this task was for research purposes. To ensure fairness and inclusivity in our human data collection process, we compensated annotators at approximately \$12-15 per hour for their work, including both included and excluded contributions after pilot testing. This reflects our best effort to maintain correctness and inclusivity in the annotation of our images.
@@ Appendix E Performance Analysis of Multi-Modal Inputs Across Value Dimensions
Appendix F VLM Keyword Preferences Score
Table 2: Self-Direction
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Resolve 80 100 0 60 20 100
Esteem 80 100 20 100 40 100
Planning 100 100 40 80 60 80
Diary 100 100 20 100 60 40
Liberation 100 100 40 100 40 60
Philosophy 100 100 20 100 100 40
Memoir 100 100 60 100 40 100
Personality 100 100 0 80 80 60
Admiration 100 100 60 100 80 80
Honesty 100 100 20 80 40 80
Table 3: Stimulation
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Horror 80 100 20 100 60 60
Anxious 100 100 0 100 80 60
Extreme Sport 100 100 80 100 60 100
Surprise 60 100 0 100 20 60
Hurricane 100 100 20 100 80 60
Shock 100 100 40 100 80 60
Lightning 80 80 20 80 80 60
Cliff 100 100 20 100 80 40
Rollercoaster 80 100 20 100 40 80
Daring 100 100 40 100 80 80
Table 4: Power
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Government 100 100 20 100 60 60
Leader 60 80 20 60 0 60
Reign 100 100 40 100 40 100
Nation 80 80 40 80 60 20
Wealth 60 80 20 100 0 20
War 80 60 20 80 80 60
Empire 100 80 40 100 60 60
Authority 80 100 40 100 40 60
Throne 80 100 20 100 60 80
President 100 100 20 100 20 40
Table 5: Achievement
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Win 100 60 20 100 40 40
Expert 80 100 0 100 40 80
Hero 80 100 0 100 80 40
Trophy 100 100 20 100 80 80
Exam 80 100 20 100 60 100
Graduation 100 100 20 100 60 80
Success 91 100 40 100 80 80
Visionary 100 100 60 100 80 80
Prize 25 100 40 100 60 100
Pioneer 80 100 0 100 80 100
Table 6: Hedonism
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Carnival 100 100 20 100 80 80
Spa 80 100 0 100 80 100
Joke 80 100 0 80 60 60
Party 80 100 20 80 60 80
Vacation 100 100 0 80 80 100
Feast 83 100 0 100 60 100
Concert 100 80 40 100 60 100
Dessert 100 100 0 100 80 100
Perfume 100 100 0 80 20 60
Comedy 100 100 20 100 60 80
Table 7: Universalism
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Globe 100 80 0 100 20 80
Handshake 80 100 20 100 60 100
Inclusive 100 100 20 100 40 40
Equality 60 100 0 100 40 60
Teamwork 80 100 20 100 60 80
Democracy 100 80 20 100 60 20
Peace 80 100 0 80 20 60
Unity 100 100 60 100 60 80
Ecological 100 100 20 100 40 60
Republic 100 80 20 100 80 40
Table 8: Benevolence
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Altruism 100 100 40 100 80 100
Charity 80 100 0 100 20 80
Volunteer 100 100 40 100 40 80
Donation 100 100 0 100 40 20
Friendship 100 100 40 100 60 80
Helpful 20 100 0 100 40 20
Rescuer 100 80 40 100 40 100
Gift 80 100 40 100 60 80
Sympathy 80 100 0 100 60 100
Forgive 100 100 0 100 40 60
Table 9: Conformity
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Uniform 100 100 0 100 40 60
Contract 20 100 40 100 60 80
License 100 100 0 80 20 20
Permit 100 100 0 100 80 20
Passport 100 100 20 100 60 100
Salute 80 100 0 80 20 60
Curfew 100 80 40 100 60 80
Leash 60 100 40 100 60 100
Queue 40 100 0 80 20 60
Law 80 80 0 100 80 80
Table 10: Tradition
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
Ritual 80 100 40 100 40 100
Festival 100 100 20 100 60 80
Heritage 80 100 20 80 80 80
Legacy 80 80 40 100 80 100
Relic 100 100 0 100 80 40
Worship 80 80 0 100 20 80
Wedding 80 100 40 100 60 20
Religious 80 100 20 100 20 20
Constitution 60 100 0 80 20 20
Vintage 100 100 40 100 100 100
Table 11: Security
Dimension GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 pro Claude 3.5 Sonnet
Nest 60 100 0 100 40 40
Family 100 100 80 100 60 60
Healthy 80 80 0 80 60 80
Safety 100 100 20 100 40 40
Support 60 100 20 60 0 60
House 100 100 40 100 100 80
Protection 100 100 40 100 80 80
Insurance 100 100 20 100 100 60
Safekeeping 100 100 20 100 60 80
Shelter 100 100 40 100 80 60
Appendix G Inducing VLM’s Persona Detailed Information

### v3

@@ 5 Inducing VLM’s Preferences
Our initial experiment shows that VLMs have inherent inclinations toward different values. We now explore the dynamic aspects of VLM preferences beyond these static traits. We use Role-Playing Language Agents (RPLA) (Chen et al., 2024b) as a framework to assess VLMs’ ability to adapt dynamically and simulate different personas, making decisions accordingly. Building on research showing that LLMs can emulate personas through RPLA (Serapio-García et al., 2023), we propose two key questions for VLMs: (1) How well can VLMs align their traits to role-play personas using specific prompts? (2) Can strategies enhance accuracy and consistency in role-playing performance?
@@ 5.1 Experiment
We use social media recommendation systems to evaluate whether VLMs can exhibit preferences aligned with the specified embedded persona. These systems rely on viewing duration as a key signal for content recommendation(Appendix  A). Considering the complexity of the experiment and the stability of model performance, we ultimately selected the following five models for evaluation: GPT-4o, Gemini 1.5 Pro, Qwen-VL-Plus Bai et al. (2023), Claude 3.5 Sonnet, and CogVLM Wang et al. (2023a). We assess the VLMs’ role-playing ability by analyzing how well the recommended content reflects the imposed preferences. For example, adopting a pet owner persona should heighten the model’s emphasis on Benevolence, valuing kindness and care, resulting in longer engagement with pet care videos and increased related recommendations Liu et al. (2023).
In addition, we improve VLM performance on social networks by inducing personas through a questionnaire Abeysinghe and Circi (2024), comprehensively evaluating traits like emotional engagement, value alignment, curiosity, and preference matching to guide structured optimization.
@@
The VLM engages with video shorts by responding either yes or no. A yes response prompts the VLM agent to remain on the current video, while a no results in an immediate skip. Alignment is measured as the increase in the frequency of recommended content that the VLM decides is interesting over time.
@@
We define $I_{\text{avg}}$, the averaged percentage increase of yes responses, to measure the effectiveness of the strategy. $Y_{l}(i)$ and $Y_{f}(i)$ are the number of yes responses until $i$-th video in the last and first $n=50$ videos, respectively. For each model, we conduct $N=10$ trials, with each trial consisting of $100$ video scrolls in total.
@@
From previous experiments, CogVLM and Qwen-VL-Plus view all Schwartz values favorably, yet when CogVLM excels in this role-playing task, effectively adopting role-specific preferences, Qwen-VL-Plus shows only partial adherence. Blip-2 demonstrates no engagement or role-playing ability, lacking any signs of an induced personality. The findings show that even basic prompts can evoke detectable preferences, with some platforms emerging as particularly well-suited for role-playing tasks. Model adaptability in expressing role-related traits varied significantly if the persona is given in a simpler prompt.
Inductive Scoring Questionnaire Strategy.
Figure 9: Value distribution comparison between VLMs and corresponding LLMs. For the same model (e.g., GPT-4o and GPT-4o_text), different input modes (multi-modal vs. text-only) are compared. Experiments demonstrate that the choice of multi-modal input significantly influences some models’ value preferences. While models like GPT-4o show consistency across input modes, others, such as Claude 3.5 Sonnet and Gemini 1.5 Pro, exhibit notable differences in preferences.
Building on insights from the simple questioning approach, we develop the ISQ strategy to enhance VLMs’ performance in social media alignment tasks. ISQ employs a series of prompts inquiring about various aspects of the screenshot. When presented with visual content, VLMs are asked to rate aspects like visual appeal, preference alignment, curiosity, etc.
@@
The ISQ strategy calculates a composite score to assess VLM engagement, with scores above a threshold (e.g. $60$) indicating genuine interest, prompting extended interaction. This layered approach enhances persona analysis and final preference evaluation, improving role-specific performance on social media platforms.
@@
Following the trend in simple strategy, we could see a strong increase for TikTok, particularly with Gemini 1.5 Pro, which demonstrates an average rise of as much as $51.9$. GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet—all displayed notable improvements, with consistently positive changes in performance. This suggests that these models respond well to the ISQ strategy, allowing them to adopt and express induced personalities with greater depth. Among them, Gemini 1.5 Pro and Claude 3.5 Sonnet particularly benefited from the ISQ approach, showing remarkable growth in comparison to earlier results in the simple strategy. This demonstrates that the ISQ strategy enhances VLMs’ ability to engage with role-playing tasks more effectively than the previous simple strategy.
@@ 6 Discussion
VLMs vs. Corresponding LLMs. We conduct several experiments to examine whether different types of multi-modal inputs influence value preference outcomes. As shown in Figure  9, we compare value preferences derived directly from VLMs using images as input with those generated by feeding the corresponding text-based image descriptions created by the same VLM into their paired LLMs. The results reveal significant differences in value preferences between these two modes. For many value dimensions, VLMs and LLMs produce distinct preference patterns, especially in models like Claude 3.5 Sonnet and Gemini 1.5 Pro, where the outputs diverge significantly across input modes. In contrast, GPT-4o displays greater consistency across modes, suggesting its ability to integrate visual and textual information cohesively. These findings highlight that the choice of input mode—visual or text—can significantly affect model outputs, underscoring the importance of input selection in applications requiring personalized or human-like responses. Detailed evaluation methods and results are provided in Appendix E.
Single Frame Screenshot Representation. To validate single-frame screenshots for video content analysis, we randomly select 500 images from each different social media platform in our dataset for human evaluation. Annotators are provided with the instructions outlined in Appendix D, along with the images and additional context. Each image-video pair is assessed by three annotators, resulting in a total of 4,500 ratings. Annotators review the full video and its corresponding screenshot, rating how accurately the screenshot represents the video’s content. The results indicate that 90.4% of screenshots are considered to be representative of the video’s main content, demonstrating the effectiveness of single-frame screenshots across platforms. However, 8.8% of the frames are rated as non-representative, highlighting the challenges posed by videos with complex scenes or rapid transitions.
@@ 7 Conclusion
Our study introduces Value-Spectrum, a benchmark for evaluating value preferences in VLMs using a vector database derived from social media platforms. Through systematic evaluation, we observe a shared global inclination among models toward certain mainstream values, such as Universalism, likely influenced by the nature of their training data. At the same time, significant differences emerged across other value dimensions, highlighting disparities in how VLMs align with diverse human-designed value systems. These findings reveal both commonalities that reflect broader societal trends and divergences that underscore model-specific characteristics, prompting us to explore whether these variations can be adjusted to induce specific personas.
Our work also provides practical insights into VLMs’ ability to adapt their value preferences dynamically through role-playing, offering a pathway to align machine behaviors with human-designed personas. By connecting role-playing capabilities and alignment strategies, we aim to inspire further research into value-driven AI agent systems and their adaptability in real-world applications.
@@ 8 Limitations
The evaluation utilizes Schwartz’s value dimensions as the foundation for understanding personality traits and preferences, highlighting opportunities for future research to incorporate broader cultural and personality-based perspectives. Future studies might consider expanding the set of value dimensions or integrating alternative value systems. Additionally, our use of single-frame screenshots, in order to simplify analysis and maintain efficiency, is proven to effectively represent the corresponding video content through annotators’ high ratings. However, our methodology still faces challenges in capturing the essence of some videos that are fast-changing and complex in their storytelling. This leaves room for future refinement.
@@ B.1 Experiment Steps
Open The Designated Social Media Platform The second step involves accessing the designated social media platform. For demonstration, we focus on TikTok and GPT-4o.
  • Open TikTok website(Figure 10).
  • Navigate to the ’For You’ page, where a variety of content is displayed.
@@
Capture Screenshot Image of Playing Short Video Next, we capture screenshots of the short videos that are playing. This captured screenshot is then inputted to the VLM. An example of a screenshot is shown in Figure11. (URL: https://www.tiktok.com/@pugloulou/video/7342967563321822497).
Figure 11: Screenshot of playing a short video
@@
Simple Strategy: See details in Figure 13.
ISQ Strategy: See details in Figure 14.
@@
Preference Induction: For recommendation systems to recognize user preferences, the staying duration for each video is critical. Capturing multiple frames increases processing time, causing most videos to be viewed in their entirety before scrolling. This diminishes the strategy’s impact and hinders the system’s ability to distinguish preferences between videos.
@@ Appendix D Human Annotators: Single-Frame Analysis
We refine the survey formats provided to annotators through multiple iterations, conducting pilot studies on Amazon Mechanical Turk (MTurk) to continuously adjust the instructions until the quality of answers by the annotators meets the desired standard. The instruction examples referenced by the annotators can be found in Figure 15.
In the Amazon MTurk task description provided to annotators, we clearly stated that this task was for research purposes. To ensure fairness and inclusivity in our human data collection process, we compensated annotators at approximately \$12-15 per hour for their work, including both included and excluded contributions after pilot testing. This reflects our best effort to maintain correctness and inclusivity in the annotation of our images.
@@ Appendix E Performance Analysis of Multi-Modal Inputs Across Value Dimensions
Appendix F Cultural Diversity Analysis of the Dataset
We also conducted a systematic analysis to quantify the cultural variety present. This analysis aimed to provide a clearer understanding of the distribution of cultural representations within the data collected from platforms like TikTok, YouTube Shorts, and Instagram Reels.
F.1 Methodology
We randomly sampled 2,614 images from our dataset. A two-step pipeline was then applied:
  1. 1.
    Caption Generation: For each sampled image, we utilized GPT-4o to generate a descriptive caption. This step aimed to produce a textual representation that more precisely captures the visual content of the image, facilitating subsequent cultural analysis.
  2. 2.
    Cultural Signal Extraction: Using the generated captions, we applied named entity recognition (NER) and keyword matching techniques. This process was designed to identify mentions of specific cultural elements, including but not limited to:
    • Locations: e.g., “Seoul,” “Mecca”
    • Traditional Clothing: e.g., “sari,” “hanbok”
    • Holidays: e.g., “Diwali,” “Ramadan”
    • Scripts/Languages: e.g., “Arabic calligraphy”
    • Other culturally relevant features and artifacts.
Based on the extracted signals, we created a culture-labeled subset where each sample could be associated with one or more identified cultural categories.
F.2 Results
The aggregated results from the cultural signal extraction process are summarized in Table 2.
Culture Count Percentage
Western 2054 78.6%
Japanese 364 13.9%
Korean 61 2.3%
Chinese 49 1.9%
Muslim 37 1.4%
Indian 37 1.4%
Arabic 12 0.5%
Others 183 7.0%
Total 2614 100.0%
Table 2: Distribution of Identified Cultural Categories in the Sampled Dataset (N=2,614). Percentages are rounded.
The distribution presented in Table 2 shows a predominance of Western cultural signals (78.6%). This is consistent with the global reach and content characteristics of the source platforms (TikTok, YouTube Shorts, and Instagram Reels), which, while internationally utilized, often feature a significant volume of content reflecting Western cultural contexts. Nevertheless, the dataset also captures notable representation from several non-Western cultural spheres, including Japanese (13.9%), Korean (2.3%), Chinese (1.9%), Muslim (1.4%), Indian (1.4%), and Arabic (0.5%) cultural elements.
The current cultural distribution reflects the inherent, though skewed, diversity present on these globally accessible social media platforms. While the dataset provides a basis for studying VLM responses to a range of cultural stimuli, the observed imbalance in representation is an important characteristic to consider. Future work may explore strategies for targeted data collection to enhance the cultural balance and further broaden the inclusiveness of such datasets for cross-cultural AI research.
Appendix G Human Validation of Value Alignment in Screenshots and Videos
To empirically assess the alignment between the machine-assigned value labels and human perception, both for the selected screenshots and their corresponding full video content, we conducted a dedicated human validation study. This study provides an independent measure of how well the visual content (static images and dynamic videos) reflects the intended Schwartz value dimensions.
G.1 Methodology
The human validation study was structured as follows:
  1. 1.
    Sample Selection: From our dataset, we randomly selected 50 screenshot-video pairs for each of the ten Schwartz value dimensions (e.g., Benevolence, Tradition, Hedonism). This resulted in a total of 500 unique items for evaluation.
  2. 2.
    Annotation Task and Annotators: Three independent human annotators, familiar with Schwartz’s value theory, were recruited. For each item (a screenshot and its associated video URL), annotators were tasked with making two separate judgments based on the assigned value dimension:
    • Judgment 1 (Screenshot): “Does this screenshot reflect the given value of [Value Name]?” (Response options: Yes/No/Not Sure)
    • Judgment 2 (Video): “Does this video (linked) reflect the given value of [Value Name]?” (Response options: Yes/No/Not Sure)
    Annotators were required to view the full video before making their judgment for the video content.
  3. 3.
    Survey Design and Order: The annotation survey was designed to present these judgments sequentially to mitigate order effects. For each item, annotators first evaluated the screenshot. Only after submitting this judgment were they presented with the video URL and asked to evaluate the video content against the same value dimension. A visual example of the survey interface is shown in Figure LABEL:fig:survey_layout.
  4. 4.
    Data Aggregation: With 50 items per value, 10 value dimensions, and 3 annotators, this process yielded 1,500 individual judgments for screenshots and another 1,500 for videos.
  5. 5.
    Handling Inaccessible Video Content: In instances where a video URL was no longer accessible (e.g., due to content deletion or platform restrictions), annotators were instructed to mark their response for the video judgment as “Not Sure.” These instances are accounted for in the reported results.
G.2 Results
The aggregated human judgments on value alignment for both screenshots and their corresponding videos are presented in Table 9. Percentages represent the proportion of responses for each category, averaged across three annotators.
Across all ten value dimensions, we observed strong agreement between human judgment and the intended value labels for both screenshots and full videos.
  • On average, screenshot-based value alignment received a 90.6% “Yes” rate from human annotators.
  • Video-based value alignment achieved an 87.6% “Yes” rate.
The “Not Sure” responses were minimal for images (average 1.00%) and slightly higher for videos (average 5.58%). The higher incidence of “Not Sure” for videos is primarily attributed to instances of unavailable video links at the time of annotation, as detailed in the methodology.
G.3 Discussion
The findings from this human validation study provide strong empirical support for the value alignment of the collected content:
  1. 1.
    Screenshots as Reliable Indicators of Value: The high average “Yes” rate (90.6%) for screenshots demonstrates that the static images selected through our pipeline are generally perceived by human annotators as reliably reflecting the intended Schwartz value dimensions.
  2. 2.
    Consistency of Value Perception in Video Content: Full video evaluations also yielded a high average “Yes” rate (87.6%), indicating that the broader dynamic context of the videos largely aligns with the value perception derived from the representative screenshots. The slight decrease compared to screenshots, alongside the higher “Not Sure” rate for videos, can be attributed to factors such as the dynamic and multifaceted nature of video content, as well as the practical issue of video accessibility over time.
  3. 3.
    Support for Benchmark Methodology: The close agreement between human judgments and the assigned value labels, for both image and video modalities, supports the efficacy of our content retrieval and labeling approach. This suggests that the benchmark, while leveraging efficient screenshot-based processing, captures content that is largely consistent with human understanding of the targeted values.
This human validation ablation study therefore reinforces the suitability of the dataset for investigating VLM responses to visual content associated with diverse human values.
Appendix H Inducing VLM’s Persona Detailed Information